OpenStudy (loser66):

Show that \(X_n= e^{sin(5n)}\) has a convergent subsequence. Please, help

OpenStudy (loser66):

@nerdguy2535

myininaya (myininaya):

well from wolfram i get that the limit does not exist as n goes to infinity but it does say \[\frac{1}{e} <e^{\sin(5n)}<e \]

myininaya (myininaya):

which means we should be able to find a sub-sequence that converges to 1/e

myininaya (myininaya):

or even e

myininaya (myininaya):

i think

myininaya (myininaya):

i don't know if that helps or not

myininaya (myininaya):

http://www.wolframalpha.com/input/?i=y%3De%5E%28sin%285x%29%29 looks like the graph is bounded so the sequence is bounded and there is a theorem you can use here called the bolzano-weiestrass theorem

OpenStudy (loser66):

Thank you very much.

OpenStudy (loser66):

i got it. since \(-1\leq sin(5x) \leq 1\)--> \(e^{sin(5x)}\) is bounded as: \(\dfrac{1}{e} \leq e^{sin(5x)}\leq e\) which show it is bounded --> it has a convergent subsequence.